A sum of money is placed at Simple Interest for 3 years at 10% p.a. and then the amount is invested for 2 years at the same rate of compound interest. If the total amount of 5 years became Rs 471,900, what was the sum?

Grade:10

2 Answers

Arun
4 years ago
 Dear student
 
let money = x
 
x * 10/100* 3 = 0.3 x
 
Now 1.3 x ( 1+ 10/100)^2 = 471900
 
1.3 x * 121/100 = 471900
 
x = 47190000 / 121 * 1.3
x = 300000
 
hence sum = 300000 ( 3 Lacs)
Vikas TU
4 years ago
Hiii 
The short method of solving this questions is :
471900÷1.573=300000
so answer will be 3lakhs
Good luck 
Cheers